OpenStudy (aonz):

If logb(a) = p and c=a^2 find in terms of p logc(b) Halp PLEASE Medal + FAN

OpenStudy (aonz):

@matricked can you please help again?

jimthompson5910 (jim_thompson5910):

Hint: logb(a) = p log(a)/log(b) = p ------------------------------ c = a^2 loga(c) = 2 log(c)/log(a) = 2 ------------------------------ See what you can do with the equations log(a)/log(b) = p log(c)/log(a) = 2

OpenStudy (anonymous):

already u have with u the value of logb(c) use loga (b) = 1/logb (a)

OpenStudy (aonz):

im lost

jimthompson5910 (jim_thompson5910):

do you see how I got the equations log(a)/log(b) = p log(c)/log(a) = 2

OpenStudy (aonz):

change base formula right?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

now multiply the two equations together

OpenStudy (aonz):

log(c)/log(a) = 2 How did you get this?

jimthompson5910 (jim_thompson5910):

more like log(c)/log(b) = 2p

jimthompson5910 (jim_thompson5910):

notice how the log(a) terms will cancel

jimthompson5910 (jim_thompson5910):

now flip both sides to get log(b)/log(c) = 1/(2p) then use the change of base rule logc (b) = 1/(2p)

OpenStudy (aonz):

i finally understand this now thanks

jimthompson5910 (jim_thompson5910):

np